Part A Part complete Calculate the voltage required to accelerate a beam of protons initially at...

Part A

Part complete

Calculate the voltage required to accelerate a beam of protons initially at rest if they have a kinetic energy of 2.9 eV .

Part B

Calculate their speed if they have a kinetic energy of 2.9 eV .

Part C

Calculate the voltage required to accelerate a beam of protons initially at rest if they have a kinetic energy of 3.3 keV .

Part D

Calculate their speed if they have a kinetic energy of 3.3 keV

Part E

Calculate the voltage required to accelerate a beam of protons initially at rest if they have a kinetic energy of 3.0×10−16 J .

Part F

Calculate their speed if they have a kinetic energy of 3.0×10−16 J .
Answer #1

a)

w = qV

2.9 e = eV

V = 2.9 V

========

b)

K = 0.5 m v^2

2.9* 1.6*10^-19 = 0.5* 1.6 *10^-27* v^2

v = 23573 m/s

=======

c)

again,

K = e V = 3300 e

V = 3300 V

=======

d)

0.5* m v^2 = 3.3*1000* e

0.5*1.67* 10^-27 * v^2 = 3300*1.6*10^-19

v = 795195 m/s

=========

e)

k = q V

3*10^-16 = 1.6*10^-19* V

V = 1875 V

=========

f)

3*10^-16 = 0.5* 1.67*10^-27 * v^2

v = 599401 m/s

=========

Comment in case any doubt, will reply for sure.. Goodluck